OpenStudy (anonymous):

how can i find the area of the region enclosed by these curves and line: 4x^2+y=4 and x^4-7=1

OpenStudy (anonymous):

actually these are just curves

OpenStudy (anonymous):

should i first try and solve for y

OpenStudy (anonymous):

r u sure the 2nd equation is right? r u missing a y in there somewhere i mean?

OpenStudy (anonymous):

yeah and also draw them to see which id the upper/lower function

OpenStudy (anonymous):

*is

OpenStudy (anonymous):

the second funtion is x^4-y=1

OpenStudy (anonymous):

ok as nadeem said u need to kno the upper and lower function. also u need to find where they intersect

OpenStudy (anonymous):

so i solved for y and got: y=4-4x^2 and y=-1+x^4

OpenStudy (anonymous):

huh, the upper curve is y=4-4x^2

OpenStudy (anonymous):

good now find where the intersect by setting the equations equal to each other and solve. so solve: 4-4x^2=-1+x^4

OpenStudy (anonymous):

If I'm not mistaken 4-4x^2 is you upper function

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

they are intesecting parabolas

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

what are the intersection points

OpenStudy (anonymous):

The intersecting points are -1 and 1

OpenStudy (anonymous):

i am trying to solve byy setting them equal to each other but.....

OpenStudy (anonymous):

4-4x^2=-1+x^4 x^4+4x^2-5=0 (x^2+5)(x^2-1) (x^2+5)(x+1)(x-1) \[x= \pm1\]

OpenStudy (anonymous):

you should get : 4(1-x^2)=x^4-1 the LHS will be zero when x=+1,-1 which is the same as the RHS bc of the even powers

OpenStudy (anonymous):

so limits of integration are -1 to 1

OpenStudy (anonymous):

wait is this symmetry??

OpenStudy (anonymous):

you should realize something right of the back, sine both functions are symmetric with the y axis, all you have to do is take the limits on integration fron [0, 1] and multiply the area by 2

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

yeah, very nice

OpenStudy (anonymous):

makes life much easier for me

OpenStudy (anonymous):

lol... sure does

OpenStudy (anonymous):

so i got the integration from 0 to 1 of 5-4x^2-x^4

OpenStudy (anonymous):

and double that:)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yyup or u can integrate it from -1 to 1 which ever u want

OpenStudy (anonymous):

hahaha.... you didn't feel like going through latex

myininaya (myininaya):

5+5-4/3-4/3-1/5-1/5 =10-8/3-2/5 =104/15

OpenStudy (anonymous):

104/15

OpenStudy (anonymous):

wow, thanks for that attachtment myiniaya, that was alot of work

OpenStudy (anonymous):

and thanks again nadeem

OpenStudy (anonymous):

no problem

myininaya (myininaya):

np

myininaya (myininaya):

don't worry about what i just typed

myininaya (myininaya):

lol

OpenStudy (anonymous):

lol